### v2

@@ 1 Introduction
While chaos in the classical domain is a well-defined phenomenon, particularly through the understanding of initial perturbations in phase space trajectories, its extension to the quantum domain presents formidable challenges. This is largely due to the breakdown of the notion of well-defined trajectories in phase space beyond the semiclassical regime, necessitating an entirely different formulation. As such, quantum chaos has traditionally been studied through correlations among the eigenvalues of the Hamiltonian, an approach pioneered by the seminal works of Wigner, Dyson, and Mehta Wigner1 ; Dyson1962a ; mehta1960statistical . These efforts sparked a wave of research, laying the foundation for studying chaotic quantum systems through statistical measures berry77 ; berry85 ; Guhr:1997ve ; Oganesyan:2007wpd ; Atas2013distribution ; Brezin1997 ; Cotler:2016fpe ; Gharibyan:2018jrp .
@@
Ever since its proposal as a toy model to explain the behavior of energy levels in atomic spectra RPmodel , the RP model has garnered significant interest as an exemplary framework for studying many-body localization, ergodicity, and quantum chaos BREZIN1996697 ; Altland_RP ; BogomolnyRP ; kravtsov2015random . Experimental findings Zhang:2023rpm have demonstrated its suitability in describing spectral transitions from uncorrelated to correlated spectra Guhr_transition ; Guhr2 ; KunzRP . Various adaptations of the RP model have been explored in the literature, including different Dyson symmetry classes Cadez:2024mjc , generalized versions including circular, log-normal, and Lévy ensembles grp1 ; PoisRP ; Buijsman:2021xbi ; lognormalRP0 ; lognormalRP ; LevyRP ; Safonova:2024ffl ; LevyRP2 , and non-Hermitian generalizations nonHRP . The non-Hermitian case is particularly illuminating, as the presence or absence of the fractal phase depends on whether the diagonal potential is real or complex. Several tools have been employed to investigate this model, such as perturbation theory Altland_RP , supersymmetry techniques Guhr_transition ; Guhr2 ; Truong:2016wcq , entropic quantities Pino_2019 ; KhaymovichKL , adiabatic gauge potentials Cadez:2024mjc , Krylov space methods Bhattacharjee:2024yxj , the replica approach Venturelli:2022hka , and the geometry of Hilbert space Sharipov:2024lah .
@@
In recent years, the theory of free probability has emerged as a powerful framework for re-examining quantum chaos from a novel perspective. Originally developed by Voiculescu voiculescu1986addition ; Voiculescu1991 ; Voiculescu1992 in the context of group isomorphisms and operator algebras, free probability theory was later expanded through its close connections with random matrix theory speicher2019lecturenotesfreeprobability ; mingo ; nica2006lectures ; Rao_Edelman_2006 ; edelmanfree , and began gaining traction in physics literature during the 1990s. Early interest was sparked by the works, among others, Gopakumar:1994iq ; Douglas:1994zu ; Zee:1996qu , who explored its relevance to large $N$ gauge theories and matrix models. More recently, the connection between free probability and generalized versions of the ETH has renewed interest in its role in understanding thermalization in quantum systems Pappalardi:2022aaz ; Pappalardi:2023nsj ; Fritzsch:2024qjn ; Wang:2025mzl ; Vallini:2024bwp ; Fava:2023pac . This resurgence has led to a range of developments, including the application of free probabilistic techniques in quantum gravity, where they offer novel insights into entanglement structure and black hole dynamics, particularly in extensions of the “island” paradigm Penington:2019kki ; Wang:2022ots . Notably, recent works Chen:2024zfj ; Camargo:2025zxr have identified a link between operator statistics and the emergence of asymptotic freeness, while other studies Jindal:2024zcg have suggested freeness as a potential missing link between different avatars of quantum chaos.
@@
The present manuscript is structured as follows. Section 2 begins with a brief description of free probability, and we also summarize some known results related to the eigenvalue statistics of the sum of two free operators. We introduce the RP model and its spectral statistics in Sec. 3. It is followed by Sec. 4, where the derivation of the DOS of the RP model using the tools of free probability is presented. In section 5 we compute various $2n$-point correlation functions, which is followed by an in-depth analysis of operator statistics in Sec. 6. The notion of free time is introduced, and its various properties are quantified. Section 7 concludes the paper with a summary and some open-ended questions.
@@ 2 Elements of Free Probability theory
Free probability theory is the branch of mathematics that studies non-commutative probability spaces. In this framework, the classical notion of independence between random variables is replaced by the notion of free independence or freeness. In this section, we provide a brief review of free probability theory and how to use it to characterize chaotic dynamics.
@@
Given a non-commutative probability space $(\mathcal{A},\varphi)$, we say that two elements $a,b\in\mathcal{A}$ are free (or freely independent) if the mixed moments of centered polynomials in $a$ and $b$ vanish whenever the variables alternate and each term is individually centered. More precisely, for all $n\geq 1$ and for any polynomials $P_{1},Q_{1},\dots,P_{n},Q_{n}$ in one variable, the following condition holds when $a$ and $b$ are free  mingo ; speicher2019lecturenotesfreeprobability ; nica2006lectures
$\left\langle(P_{1}(a)-\langle P_{1}(a)\rangle)(Q_{1}(b)-\langle Q_{1}(b)\rangle)\cdots(P_{n}(a)-\langle P_{n}(a)\rangle)(Q_{n}(b)-\langle Q_{n}(b)\rangle)\right\rangle=0\,.$ (1)
@@
In the simplest case where all polynomials are linear, i.e., $P(x)=Q(x)=x$, this definition yields:
@@
With this definition of the free cumulants, one can express the condition of freeness between two variables $A$ and $B$ as the statement that all the mixed free cumulants of these two variables are zero mingo ; nica2006lectures , i.e., $\kappa_{n}(C_{1},\cdots C_{n})=0$ for $n\geq 2$, $C_{i}\in\{A,B\}$ for all $i$, and $i$ and $j$ such that $C_{i}=A,C_{j}=B$. One straightforward conclusion of this fact, along with the relation between the moments and free cumulants, is that all non-crossing moments containing free variables get factorized in terms of moments of the individual variables.
@@
A diagrammatic technique was used in Zee:1996qu to derive the additivity property (9) of the $R$-transformation for the addition of two independent large Gaussian random matrices. By noting that the Cuachy transformation $G(z)$ of the density of eigenvalues (i.e., the Green function), can be written in terms of the one-particle irreducible self-energy $\Sigma(z)$ as a Dyson-Schwinger equation $G(z)=1/(z-\Sigma(z))$, we see that the $R$-transformation is just the self-energy,
@@ 2.1 Summary of Known Results on Free Additive Convolution
In this subsection, we present some known results for the free additive convolution of free operators. We refer to Camargo:2025zxr for a detailed derivation of these results.
@@
Figure 1: The operator eigenvalue distributions predicted by the free probability for $s=1/2$ (left) $s=1$ (middle) and $s=3/2$ (right). The spin $1/2$ and the spin $1$ distribution is given by the arcsine distibution (12) and distribution (14) respectively, while the distrbution for spin $3/2$ can be evaluated using the subordination method.
@@
For operators with spin larger than unity, it becomes increasingly challenging to obtain analytic formulae for the free additive convolution of the sum of operators, due to the fact that one needs to solve order-$s$ algebraic equations. Fortunately, one can obtain such distributions numerically using subordination function methods; see, for instance speicher2019lecturenotesfreeprobability . Such methods were applied in Camargo:2025zxr for the derivation of the free additive convolution of spin-$3/2$ and spin-$2$ operators. Figure 1 shows the respective distribution for spin-$1/2$, spin-$1$, and spin-$3/2$. We review the subordination function method in Appendix B.
@@ 2.2 Asymptotic Freeness Produced by Chaotic Dynamics
Freeness arises only in infinite-dimensional Hilbert spaces. When operators $P$ and $Q$ are represented as $N\times N$ matrices, freeness can emerge only in the limit $N\to\infty$, a regime associated with asymptotic freeness. For finite $N$, cumulants never vanish completely and retain residual values at late times. However, under chaotic dynamics – i.e., when the Hamiltonian exhibits features akin to random matrix theory – the magnitude of these residual cumulants decreases with increasing $N$, suggesting their vanishing in the $N\to\infty$ limit. In this sense, the finite-$N$ system “knows” about its large-$N$ counterpart.
@@ 2.3 Relation with $2n$-point functions
[…]
When one of these operators is a time-evolved Heisenberg operator (as is the case we consider below in section 6), the moments $\langle(P+Q(t)^{k}\rangle$ have expansions in terms of the OTOCs of different orders. [^5] We thank Koji Hashimoto for discussions on this point. To see this explicitly, consider two spin-$1/2$ operators and the following second-order moment:
@@
where the normalization $\langle P^{2k}\rangle=\langle Q^{2k}\rangle=1$ has been applied throughout, and the operators are chosen such that their one-point function vanishes: $\langle P\rangle=\langle Q(t)\rangle=0$. In other words, we choose the traceless operators or the traceless part of the operators by appropriately subtracting the traces. For the random matrix ensembles, the average is taken over the ensembles of the Hamiltonians.
@@
As expected, these moments correspond directly to those of the free probability prediction for the eigenvalue distribution of the operator $P+Q(t)$ at late times. For example, it is easy to see that the moments of the arcsine distribution in (12) are given by the same binomial coefficients:
$\int_{-2}^{2}\mathrm{d}\lambda~{}\frac{\lambda^{2n}}{\pi\sqrt{4-\lambda^{2}}}=\binom{2n}{n}\,.$ (19)
Thus, for the sum of two spin-$1/2$ operators, the eigenvalue distribution converges to the arcsine law predicted by the free probability. Extensions to higher spins follow naturally.
@@ 3.1 Hamiltonian and the spectral statistics
Here, $A$ is an $N\times N$ random diagonal matrix, chosen from independent and identically distributed (iid) elements from a normal distribution with zero mean and variance $\sigma^{2}_{A}=4/N$. On the other hand, $B$ belongs to the Gaussian Orthogonal Ensemble (GOE) with zero mean and variance $\sigma^{2}_{B}=2/N$. However, the Hamiltonian can be chosen from any three symmetry classes of Dyson ensembles Cadez:2024mjc . The Hamiltonian shows rich phase structure across different parametric regimes: the ergodic regime for $0\leq\gamma\leq 1$, non-ergodic and extended fractal phase for $21<\gamma<$ 2, and localized phase for $\gamma>2$. The two transitions, namely the ergodic to fractal phase at $\gamma=1$ and the fractal phase to localized phase at $\gamma=2$, can be found analytically at the large $N$ limit. The DOS transitions from a semicircle distribution supported by the interval $[-2,2]$ for $\gamma=0$ to the normal distribution with zero mean and variance $4/N$ for large $\gamma$.
The distributions of nearest-level spacings and their ratios constitute interesting diagnostics of chaotic Hamiltonians. Denoting the eigenvalues of the $N\times N$ Hamiltonian as $\lambda_{n}$ with $n=1,\cdots,N$, the nearest level spacings are given by the difference between the nearest eigenvalues as $\xi_{k}=\lambda_{k+1}-\lambda_{k}$, with $k=1,\cdots,N-1$. For random matrix theory, such distributions follow universal statistics, given by the Wigner–Dyson distribution Haake:2010fgh :
@@
[…]
@@
[…]
@@
For different system sizes, a clear crossover at $\gamma_{c}=2$ is observed, which becomes more prominent after collapsing the data into a single curve. The conclusion drawn is that while the nearest level spacing statistics effectively mark the transition from fractal to localized phases, it fails to adequately capture the transition from ergodic to fractal phases at $\gamma=1$. It is noteworthy that such observations have already been reported in the literature Pino_2019 ; PoisRP ; Buijsman:2023ips ; Bhattacharjee:2024yxj , and efforts have been made to identify the ergodic to fractal transitions using various probes.
@@
[…]

### v3

@@ 1 Introduction
While chaos in the classical domain is a well-defined phenomenon, particularly through the understanding of initial perturbations in phase space trajectories, its extension to the quantum domain presents formidable challenges. This is largely due to the breakdown of the notion of well-defined trajectories in phase space beyond the semiclassical regime, necessitating an entirely different formulation. As such, quantum chaos has traditionally been studied through correlations among the eigenvalues of the Hamiltonian, an approach pioneered by the seminal works of Wigner, Dyson, and Mehta Wigner1 ; Dyson1962a ; mehta1960statistical . These efforts sparked a wave of research, laying the foundation for studying chaotic quantum systems through statistical measures berry77 ; Guhr:1997ve ; Oganesyan:2007wpd ; Atas_rvalue ; Brezin1997 ; Cotler:2016fpe ; Gharibyan:2018jrp .
@@
Ever since its proposal as a toy model to explain the behavior of energy levels in atomic spectra RPmodel , the RP model has garnered significant interest as an exemplary framework for studying many-body localization, ergodicity, and quantum chaos BREZIN1996697 ; Altland_RP ; BogomolnyRP ; kravtsov2015random . Experimental findings Zhang:2023rpm have demonstrated its suitability in describing spectral transitions from uncorrelated to correlated spectra Guhr_transition ; Guhr2 ; KunzRP . Various adaptations of the RP model have been explored in the literature, including different Dyson symmetry classes Cadez:2024mjc , generalized versions including circular, log-normal, and Lévy ensembles grp1 ; PoisRP ; Buijsman:2021xbi ; lognormalRP0 ; lognormalRP ; Monthus_2017 ; LevyRP ; Safonova:2024ffl ; LevyRP2 ; kutlinLevy , and non-Hermitian generalizations nonHRP exhibiting various localization phenomena madhumitaRP1 ; madhumitaRP2 . The non-Hermitian case is particularly illuminating, as the presence or absence of the fractal phase depends on whether the diagonal potential is real or complex. Several tools have been employed to investigate this model, such as perturbation theory Altland_RP , supersymmetry techniques Guhr_transition ; Guhr2 ; Truong:2016wcq , entropic quantities Pino_2019 ; lognormalRP0 , adiabatic gauge potentials Cadez:2024mjc , Krylov space methods Bhattacharjee:2024yxj , the replica approach Venturelli:2022hka , and the geometry of Hilbert space Sharipov:2024lah .
@@
In recent years, the theory of free probability has emerged as a powerful framework for re-examining quantum chaos from a novel perspective. Originally developed by Voiculescu voiculescu1986addition ; Voiculescu1991 ; Voiculescu1992 in the context of group isomorphisms and operator algebras, free probability theory was later expanded through its close connections with RMT speicher2019lecturenotesfreeprobability ; mingo ; nica2006lectures ; Rao_Edelman_2006 ; edelmanfree , and began gaining traction in physics literature during the 1990s. Early interest was sparked by the works, among others, Gopakumar:1994iq ; Douglas:1994zu ; Zee:1996qu , who explored its relevance to large $N$ gauge theories and matrix models. More recently, the connection between free probability and generalized versions of the ETH has renewed interest in its role in understanding thermalization in quantum systems Pappalardi:2022aaz ; Pappalardi:2023nsj ; Fritzsch:2024qjn ; Wang:2025mzl ; Vallini:2024bwp ; Fava:2023pac . This resurgence has led to a range of developments of free probabilistic techniques in quantum gravity, such as computing the eigenspectra Gao:2024lem of Sachdev-Ye-Kitaev models, or the entanglement spectra in extensions of the “island” paradigm Penington:2019kki ; Wang:2022ots . Notably, recent works Chen:2024zfj ; Camargo:2025zxr have identified a link between operator statistics and the emergence of asymptotic freeness, while other studies Jindal:2024zcg have suggested freeness as a potential missing link between different avatars of quantum chaos.
@@
The present manuscript is structured as follows. Sec. 2 begins with a brief description of free probability, and we also summarize some known results related to the eigenvalue statistics of the sum of two free operators. We introduce the RP model and its spectral statistics in Sec. 3. It is followed by Sec. 4, where the derivation of the DOS of the RP model using the tools of free probability is presented. In Sec. 5 we compute various $2n$-point correlation functions, which is followed by an in-depth analysis of operator statistics in Sec. 6. The notion of free time is introduced, and its various properties are quantified. Sec. 7 concludes the paper with a summary and some open-ended questions. In four appendices, we elaborate on some further important points. Specifically, in the Appendix A we compare the asymptotic freeness governed by different unitary random matrix ensembles, while in Appendix B we illustrate the role played by the delocalized eigenvectors by considering a decorrelated version of the RP random matrix ensemble. Finally, in Appendix C we review the procedure of obtaining the free convolution prediction for the sum of two spin $s$ operators using the subordination relations, and in Appendix D the operator-dependent nature of the emergence free probability prediction is examined.
@@ 2 Elements of Free Probability theory
Free probability theory is the branch of mathematics that studies non-commutative probability spaces. In this framework, the classical notion of independence between random variables is replaced by the notion of free independence or freeness nica2006lectures ; speicher2019lecturenotesfreeprobability ; mingo ; voiculescu1986addition . In this section, we provide a brief review of free probability theory and how to use it to characterize chaotic dynamics.
@@
Given a non-commutative probability space $(\mathcal{A},\varphi)$, we say that two elements $a,b\in\mathcal{A}$ are free (or freely independent) if the mixed moments of centered polynomials in $a$ and $b$ vanish whenever the variables alternate and each term is individually centered. More precisely, for all $n\geq 1$ and for any polynomials $\mathcal{P}_{1},\mathcal{Q}_{1},\dots,\mathcal{P}_{n},\mathcal{Q}_{n}$ in one variable, the following condition holds when $a$ and $b$ are free  mingo ; speicher2019lecturenotesfreeprobability ; nica2006lectures
$\left\langle(\mathcal{P}_{1}(a)-\langle\mathcal{P}_{1}(a)\rangle)(\mathcal{Q}_{1}(b)-\langle\mathcal{Q}_{1}(b)\rangle)\cdots(\mathcal{P}_{n}(a)-\langle\mathcal{P}_{n}(a)\rangle)(\mathcal{Q}_{n}(b)-\langle\mathcal{Q}_{n}(b)\rangle)\right\rangle=0\,.$ (1)
@@
In the simplest case where all polynomials are linear, i.e., $\mathcal{P}(x)=\mathcal{Q}(x)=x$, this definition yields:
@@
With this definition of the free cumulants, one can express the condition of freeness between two variables $P$ and $Q$ as the statement that all the mixed free cumulants of these two variables are zero mingo ; nica2006lectures , i.e., $\kappa_{n}(C_{1},\cdots C_{n})=0$ for $n\geq 2$, $C_{i}\in\{P,Q\}$ for all $i$, and $i$ and $j$ such that $C_{i}=P,C_{j}=Q$. One straightforward conclusion of this fact, along with the relation between the moments and free cumulants, is that all non-crossing moments containing free variables get factorized in terms of moments of the individual variables.
@@
A diagrammatic technique was used in Zee:1996qu to derive the additivity property (9) of the $R$-transformation for the addition of two independent large Gaussian random matrices. By noting that the Cuachy transformation $G(z)$ of the density of eigenvalues (i.e., the Green function), can be written in terms of the one-particle irreducible self-energy $\Sigma(z)$ as a Dyson–Schwinger equation $G(z)=1/(z-\Sigma(z))$, we see that the $R$-transformation is just the self-energy,
@@ 2.1 Summary of Known Results on Free Additive Convolution
In this subsection, we present some known results for the free additive convolution of free operators. We refer to speicher2019lecturenotesfreeprobability ; Camargo:2025zxr ; nica2006lectures for a detailed derivation of these results.
@@
Figure 1: The operator eigenvalue distributions predicted by the free probability for the addition of two free spin $s=1/2$ (left), $s=1$ (middle), and $s=3/2$ (right) operators. The spin $1/2$ and the spin 11 distribution is given by the arcsine distribution (12) and distribution (14), respectively. In contrast, the distribution for spin $3/2$ can be evaluated using the subordination method.
@@
For operators with spin larger than unity, it becomes increasingly challenging to obtain analytic formulae for the free additive convolution of the sum of operators, due to the fact that one needs to solve order-$s$ algebraic equations. Fortunately, such distributions can be obtained numerically using subordination function methods; see, for instance speicher2019lecturenotesfreeprobability . Such methods were applied in Camargo:2025zxr for the derivation of the free additive convolution of spin-$3/2$ and spin-$2$ operators. Figure 1 shows the respective distribution for spin-$1/2$, spin-11, and spin-$3/2$. We review the subordination function method in Appendix C.
@@ 2.2 Asymptotic Freeness Produced by Chaotic Dynamics
Freeness arises only in infinite-dimensional Hilbert spaces. When operators $P$ and $Q$ are represented as $N\times N$ matrices, freeness can emerge only in the limit $N\to\infty$, a regime associated with asymptotic freeness. For finite $N$, mixed cumulants never vanish completely and retain residual values at late times. However, under chaotic dynamics – i.e., when the Hamiltonian exhibits features akin to RMT – the magnitude of these residual cumulants decreases with increasing $N$, suggesting their vanishing in the $N\to\infty$ limit. In this sense, the finite-$N$ system “knows” about its large-$N$ counterpart.
@@ 2.3 Relation with $2n$-point functions
It is imperative to ask how the eigenvalue distribution of the sum of two operators is related to the higher-point correlation functions, namely the higher-order OTOCs. To determine the eigenvalue distribution of a matrix $\mathcal{O}$, [^5] We assume that the operators we consider in this paper have corresponding eigenvalue distributions with compact supports on the real line. it suffices to evaluate the traces of all powers of the operator, $\text{Tr}(\mathcal{O}^{k})$. For a random matrix $\mathcal{O}$, its ensemble-averaged eigenvalue distribution can be derived from $\langle\mathcal{O}^{k}\rangle$, where $\langle\cdot\rangle$ denotes the infinite-temperature inner product, computed as $\mathbb{E}(\mathrm{Tr}(\cdot))/N$, with $N$ representing the rank of the matrix mingo ; speicher1997free . Consequently, the ensemble-averaged eigenvalue distribution of an operator formed as $P+Q$ can be obtained from the moments $\langle(P+Q)^{k}\rangle$. Therefore, the information about the ensemble-averaged eigenvalue distribution of an operator of the form $P+Q$ can be obtained from the moments $\langle(P+Q)^{k}\rangle$, or alternatively from the mixed moments of the form $\langle P^{n_{1}}Q^{m_{1}}P^{n_{2}}Q^{m_{2}}\cdots\rangle$.
When one of these operators is a time-evolved Heisenberg operator (as is the case we consider below in Sec. 6), the moments $\langle(P+Q(t)^{k}\rangle$ have expansions in terms of the OTOCs of different orders. [^6] We thank Koji Hashimoto for discussions on this point. To see this explicitly, consider two spin-$1/2$ operators and the following second-order moment:
@@
where the normalization $\langle P^{2k}\rangle=\langle Q^{2k}(t)\rangle=1$ has been applied throughout (keeping in mind they are Pauli operators acting on qubits), and the operators are chosen such that their one-point function vanishes: $\langle P\rangle=\langle Q(t)\rangle=0$. In other words, we choose the traceless operators or the traceless part of the operators by appropriately subtracting the traces. For the random matrix ensembles, the average is taken over the ensembles of the Hamiltonians.
@@
As expected, these moments correspond directly to those of the free probability prediction for the eigenvalue distribution of the operator $P+Q(t)$ at late times. For example, it is easy to see that the moments of the arcsine distribution in (12) are given by the same binomial coefficients: [^7] The moment generating function of the arcsine distribution are given by the $I_{0}(2u)$, where $I_{n}(u)$ is the $n$-th order modified Bessel function of the first kind. It is easy to check that $m_{2n}=d^{2n}I_{0}(2u)/du^{2n}\big|_{u\rightarrow 0}$ are the binomial coefficients (19), while the odd moments vanish.
$m_{2n}=\int_{-2}^{2}\mathrm{d}\lambda~\frac{\lambda^{2n}}{\pi\sqrt{4-\lambda^{2}}}=\binom{2n}{n}\,~~~n=0,1,2,\cdots\,.$ (19)
Thus, for the sum of two spin-$1/2$ operators, the eigenvalue distribution converges to the arcsine law predicted by the free probability. Extensions to higher spins follow in a similar fashion.
@@ 3.1 Hamiltonian and the spectral statistics
Here, $A$ is an $N\times N$ random diagonal matrix, chosen from independent and identically distributed (iid) elements from a normal distribution with zero mean and variance $\sigma^{2}_{A}=4/N$. On the other hand, $B$ belongs to the Gaussian Orthogonal Ensemble (GOE) with zero mean and variance $\sigma^{2}_{B}=2/N$. However, the Hamiltonian can be chosen from any three symmetry classes of Dyson ensembles Cadez:2024mjc . The Hamiltonian shows rich phase structure across different parametric regimes: the ergodic regime for $0\leq\gamma\leq 1$, non-ergodic and extended fractal phase for $1<\gamma<2$, and localized phase for $\gamma>2$. The two transitions, namely the ergodic to fractal phase at $\gamma=1$ and the fractal phase to localized phase at $\gamma=2$, can be found analytically at the large $N$ limit. The DOS transitions from a semicircle distribution supported by the interval [2,2][-2,2] for $\gamma=0$ to the normal distribution with zero mean and variance $4/N$ for large $\gamma$.
The distributions of nearest-level spacings and their ratios constitute interesting diagnostics of chaotic Hamiltonians. Denoting the eigenvalues of the $N\times N$ Hamiltonian as $\lambda_{n}$ with $n=1,\cdots,N$, the nearest level spacings are given by the difference between the nearest eigenvalues as $\xi_{k}=\lambda_{k+1}-\lambda_{k}$, with $k=1,\cdots,N-1$. For RMT, such distributions follow universal statistics, given by the Wigner–Dyson distribution Haake:2010fgh :
@@
However, to compute the level spacing distributions, it is required to locally uniformize the DOS, such that the mean spacing becomes unity $\int_{0}^{\infty}p_{\upbeta}(\xi)\,d\xi=\int_{0}^{\infty}p_{\upbeta}(\xi)\,\xi d\xi=1$. This method, known as unfolding, often gets complicated, and there is no universal way to perform such a procedure. To avoid this, it is useful to define the distribution of nearest-level spacing ratios, known as the $r$-value statistics, which is free of such unfolding. This is defined as Oganesyan:2007wpd :
@@
Besides the distribution itself, it is often convenient to compute its mean $\langle r\rangle_{\upbeta}:=\int_{0}^{1}r\,p_{\upbeta}(r)\,dr$. For GOE, GUE, and GSE, it is given by $\langle r\rangle_{\mathrm{GOE}}\approx 0.536$, $\langle r\rangle_{\mathrm{GUE}}\approx 0.603$, and $\langle r\rangle_{\mathrm{GSE}}\approx 0.676$ respectively Atas_rvalue . For uncorrelated spectra with Poissonian statistics, it is given by $\langle r\rangle_{\mathrm{P}}\approx 0.386$.
@@
For different system sizes, a clear crossover at $\gamma_{c}=2$ is observed, which becomes more prominent after collapsing the data into a single curve. The conclusion drawn is that while the nearest level spacing statistics effectively mark the transition from fractal to localized phases, they fail to adequately capture the transition from ergodic to fractal phases at $\gamma=1$. It is noteworthy that such observations have already been reported in the literature Pino_2019 ; PoisRP ; Buijsman:2023ips ; Bhattacharjee:2024yxj , and efforts have been made to identify the ergodic to fractal transitions using various probes.
@@
Figure 2: (a) The level spacing distribution $p(\xi)$ in the ergodic ($\gamma=0)$, fractal ($\gamma=1.5$), and localized ($\gamma=5$) regimes. The solid blue and red lines denote the GOE and the Poisson distribution. The system size is $N=2^{10}$, with a $10^{4}$ ensemble of Hamiltonians taken. For the spacing, only the bulk of the eigenvalue spectrum ($20\%$ around each side of the mid-spectrum) is considered. (b) The $r$-ratio of the RP model across different phases for different sizes of the matrices $N=2^{8}\,(5\times 10^{4}),\,2^{10}\,(10^{4})$ and $2^{12}\,(10^{2})$. The ensemble averages are shown in parentheses. The inset shows the data collapse, which clearly marks the fractal to the localized transition at $\gamma_{c}=2$.